For the following ellipse:
(x + 2)2 + (y + 2)2 = 1
4 25

Find: a)Center :___________
b)Vertices:___________

___________
c) Foci: ___________
___________
d) length of
major axis:________
e) length of
minor axis:________
f) eccentricity:________


Answer by Boreal(15235)

Center is (-2,-2) You change the sign of each.
The vertices are along the y-axis, since that is how the ellipse is oriented with regard to the major axis. They are sqrt (25) or 5. They are at (-2,-7) and (-2,+3)
The major axis is 10 units long; the minor axis is 4 units long. It is the square root of the denominator multiplied by 2.
a^2-c^2=b^2
25-21=4
c^2=21;; c= sqrt(21) ; The foci are sqrt(21) from the center, along the y-axis, so they are at (-2, -2+sqrt(21) and (-2, -2- sqrt(21))
eccentricity is c/a, and that is sqrt (21)/5